We say that an nxn matrix is skew-symmetric if A^T=-A. Let W be the set of all 2x2 skew-symmetric matrices: W = {A in m2x2(R) l A^T=-A}.

(a) Show that W is a subspace of M2x2(R)

(b) Find a basis for W and determine dim(W).

(c) Suppose T: M2x2(R) is a linear transformation given by T(A)=A^T +A. Is T injective? Is T surjective? Why or why not? You do not need to verify that T is linear.

3. (17 points) Subspaces & Linear Transformations We say that an n x n matrix is skew-symmetric if AT = - A. Let W be the set(b) (5 points) Find a basis for W and determine dim(W).(c) (6 points) Suppose T : M2x2(R) + M2x2(R) is a linear transformation given by T(A) = AT + A. Is T injective? Is T surjecti
Answer #1

3 . 3. (a) W = 3AE {AE Max2 (IR) | AT=- Then (1) Let A and B in w. ATRA, B=-B we need to show (A-BT = AT BT = -A-(-3) A-BEW,(b) Let b A= and AEW. d :. AT А b d e comparing d=-d as-a, Q=0 »d **3 6,1 ---, [] w={[:] b b ·A= 1 3 ab basis forys wa i dimNow P= be a matrix non-zero 다. in Mexe (18) But T(P) = pt +P [146]= - is PE ker() Hence element, then ker (T) contains a non-

I think it will help you . If you have any doubts feel free to ask in the comments section .